### v3

# Cycles of Length 4 or 8 in Graphs with Diameter 2 and Minimum Degree at Least 3

Abstract
In this short note it is shown that every graph of diameter $2$ and minimum degree at least $3$ contains a cycle of length $4$ or $8$. This result contributes to the study of the Erdős–Gyárfás Conjecture [1] by confirming it for the class of diameter-$2$ graphs.
Notation and Preliminaries
All graphs considered in this note are finite, simple, and undirected.
Let $G=(V(G),E(G))$ be a graph where $V(G)$ and $E(G)$ denote the set of vertices and edges in G respectively. For a vertex $v\in V(G)$, the neighborhood of $v$ is
$N(v)=\{u\in V(G):uv\in E(G)\},$
and the degree of $v$ is $d(v)=|N(v)|$. The minimum degree of $G$ is denoted by
$\delta(G)=\min\{d(v):v\in V(G)\}.$
For vertices $u,v\in V(G)$, the distance $d(u,v)$ is the length of a shortest path joining $u$ and $v$ in $G$. The diameter of $G$ is
$\operatorname{diam}(G)=\max\{d(u,v):u,v\in V(G)\}.$
A path of length $k$ is a sequence of distinct vertices
$v_{0},v_{1},\dots,v_{k}$
such that $v_{i}v_{i+1}\in E(G)$ for all $0\leq i<k$. A cycle of length $k$ is a sequence
$v_{0},v_{1},\dots,v_{k-1},v_{0}$
in which $v_{0},\dots,v_{k-1}$ are distinct and $v_{i}v_{i+1}\in E(G)$ for all indices taken modulo $k$.
A $k$-cycle is a cycle of length $k$. A cycle is called simple if it contains no repeated vertices except for the initial and terminal vertex.
Given a cycle $C$, an edge joining two nonconsecutive vertices of $C$ is called a chord of $C$.
Throughout the paper, vertex labels $v_{1},v_{2},\dots$ are introduced as needed and are proved to be distinct. When a sequence is written as
$v_{1}-v_{2}-\cdots-v_{k},$
it means that $v_{i}v_{i+1}\in E(G)$ for all $1\leq i<k$.
All set-theoretic notation is used in its standard meaning.
Introduction
A well-known open problem of Erdős and Gyárfás asks for unavoidable cycle lengths in graphs with minimum degree at least three. In particular, they conjectured that every graph $G$ with $\delta(G)\geq 3$ contains a simple cycle whose length is a power of two. This is now commonly referred to as the Erdős–Gyárfás Conjecture. Folklore has the conjecture first appearing at a conference in (1995) (later in literature in (1997) [1]), and is listed in several open-problem compilations (e.g. West [2] and Erdős problems forums such as [3]).
Despite its simple statement, the conjecture remains open in full generality and has been verified only for restricted classes of graphs. Early progress includes results for planar and cubic claw-free graphs [4, 5]. More recently, Heckman and Krakovski proved the conjecture for $3$-connected cubic planar graphs [6], and a number of papers have established the conjecture for other hereditary or structured graph classes such as $P_{8}$-free graphs [7].
Computational searches have also shaped the current understanding. In particular, exhaustive and heuristic searches (notably by Royle and Markström) indicate that any counterexample must be relatively large, and Markström produced extremal examples illustrating how power-of-two cycles can be forced to occur only at larger lengths (e.g. $16$) in certain cubic graphs [8]. These computations suggest that minimal counterexamples, if they exist, are highly constrained.
In this note the Erdős–Gyárfás conjecture is verified for graphs of diameter $2$. The main result shows that every graph $G$ with $\operatorname{diam}(G)=2$ and $\delta(G)\geq 3$ contains a cycle of length $4$ or $8$ (see Theorem 1.1.). Thus, within the diameter–$2$ regime, the conjecture holds in its strongest possible form, guaranteeing one of the two smallest nontrivial powers of two. From the perspective of the broader conjecture, diameter $2$ graphs form a natural and widely studied class: the global constraint $\operatorname{diam}(G)=2$ forces any two nonadjacent vertices to have a common neighbor, creating a dense web of short connections. The proof exploits this interaction between a local degree lower bound and the global diameter constraint to force short power-of-two cycles.
Main Result
Theorem 1.1.: Let $G$ be a graph with diameter $2$ and minimum degree at least $3$. Then $G$ contains a cycle of length $4$ or $8$.
Proof
Assume $G$ has diameter $2$, minimum degree at least $3$, and no $4$-cycle. For a vertex $v$ in $G$, let $N(v)$ denote its neighborhood; the set of vertices adjacent to $v$.
Let $v_{1}v_{2}\in E(G)$. By the degree condition, $v_{1}$ has two neighbors other than $v_{2}$, call them $v_{3},v_{4}$; similarly, $v_{2}$ has two neighbors $v_{5},v_{6}$. Denote this subgraph of $G$ as $G^{\prime}$ (see Figure 1).
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$
Figure 1: $G^{\prime}$ - Initial edge with neighbors satisfying the degree constraint.
If $v_{3}=v_{5}$ while $v_{4}=v_{6}$, or $v_{3}=v_{6}$ while $v_{4}=v_{5}$, then a cycle of length $4$ forms immediately, namely
$v_{1}-v_{3}-v_{2}-v_{4}-v_{1}\quad\text{or}\quad v_{1}-v_{4}-v_{2}-v_{3}-v_{1}.$
Both cases contradict the assumption that $G$ contains no $4$-cycle.
Thus, without loss of generality, it suffices to prove the theorem by considering separately the cases $v_{3}=v_{5}$ and $v_{3}\neq v_{5}$.
Case 1: $v_{3}=v_{5}$
Assume $v_{3}=v_{5}$. Let $v_{4}$ be the neighbor of $v_{1}$ distinct from $v_{2},v_{3}$, and let $v_{6}$ be the neighbor of $v_{2}$ distinct from $v_{1},v_{3}$. Set $V(G^{\prime})=\{v_{1},v_{2},v_{3},v_{4},v_{6}\}$.
Claim 1.1. $v_{4}v_{6}\notin E(G)$.
Proof. If $v_{4}v_{6}\in E(G)$, then
$v_{4}-v_{1}-v_{2}-v_{6}-v_{4}$
is a $4$-cycle, contradicting the assumption that $G$ contains no $4$-cycle. $\diamond$
Since $v_{4}$ and $v_{6}$ are nonadjacent and $\operatorname{diam}(G)=2$, we have
$N(v_{4})\cap N(v_{6})\neq\varnothing.$
Choose
$v_{7}\in N(v_{4})\cap N(v_{6}).$
Claim 1.2. $v_{7}\notin\{v_{1},v_{2},v_{3}\}$.
Proof. If $v_{7}=x$ such that $x\in\{v_{1},v_{2},v_{3}\}$, then at least one of the following 44–cycles would form:
$v_{1}-v_{3}-v_{2}-v_{6}-v_{1},$
$v_{2}-v_{3}-v_{1}-v_{6}-v_{2},$
$v_{4}-v_{3}-v_{2}-v_{1}-v_{4}.$
$\diamond$
Thus, assume $v_{7}\notin V(G^{\prime})$. Then the closed walk
$v_{7}-v_{4}-v_{1}-v_{3}-v_{2}-v_{6}-v_{7}$
forms a $6$-cycle with a chord $v_{1}v_{2}$ (see Figure 2).
$v_{7}$$v_{4}$$v_{1}$$v_{2}$$v_{6}$$v_{3}$
Figure 2: 66-Cycle with a $v_{1}v_{2}$ Chord
Notice $v_{3}$ and $v_{7}$ are not adjacent, and if the edge $v_{3}v_{7}\in E(G)$, then a $4$-cycle forms by
$v_{7}-v_{3}-v_{1}-v_{4}-v_{7},$
a contradiction. Also, the diameter–$2$ condition implies
$N(v_{3})\cap N(v_{7})\neq\varnothing.$
Let $v_{8}\in N(v_{3})\cap N(v_{7})$.
Claim 1.3 $v_{8}\notin V(G^{\prime})$.
Proof.
Indeed, if $v_{8}\in V(G^{\prime})$, then $v_{8}\in\{v_{1},v_{2},v_{4},v_{6}\}$ and a 44–cycle arises in each case:
$v_{8}=v_{1}$ $\implies v_{1}-v_{2}-v_{6}-v_{7}-v_{1},$
$v_{8}=v_{2}$ $\implies v_{2}-v_{1}-v_{4}-v_{7}-v_{2},$
$v_{8}=v_{4}$ $\implies v_{4}-v_{3}-v_{2}-v_{1}-v_{4},$
$v_{8}=v_{6}$ $\implies v_{6}-v_{3}-v_{1}-v_{2}-v_{6}.$
This contradicts the assumption that $G$ contains no 44–cycle. Hence $v_{8}\notin V(G^{\prime})$
(see Figure 3). $\diamond$
$v_{7}$$v_{4}$$v_{1}$$v_{2}$$v_{6}$$v_{3}$$v_{8}$
Figure 3: $v_{8}\in N(v_{7})\cap N(v_{3})$
Since $\delta(G)\geq 3$, the vertex $v_{8}$ has a neighbor $v_{9}\notin\{v_{3},v_{7}\}$.
Claim 1.4. $v_{9}\notin V(G^{\prime})$.
Proof. Otherwise $v_{9}\in\{v_{1},v_{2},v_{4},v_{6}\}$, and each possibility yields a 44–cycle:
$v_{9}=v_{1}$ $\implies v_{1}-v_{4}-v_{7}-v_{8}-v_{1},$
$v_{9}=v_{2}$ $\implies v_{2}-v_{1}-v_{3}-v_{8}-v_{2},$
$v_{9}=v_{4}$ $\implies v_{4}-v_{8}-v_{3}-v_{1}-v_{4},$
$v_{9}=v_{6}$ $\implies v_{6}-v_{2}-v_{3}-v_{8}-v_{6}.$
$\diamond$
Thus $v_{9}\notin V(G^{\prime})$ (see Figure 4).
$v_{7}$$v_{4}$$v_{1}$$v_{2}$$v_{6}$$v_{3}$$v_{8}$$v_{9}$
Figure 4: $v_{8}v_{9}\in E(G)$
Claim 1.5. If $v_{9}$ is adjacent to $x\in\{v_{1},v_{2},v_{4},v_{6}\}$, then there exists a 44–cycle.
Proof.
Assume $v_{9}$ is adjacent to $x\in\{v_{1},v_{2},v_{4},v_{6}\}$. Thus, a 44–cycle is formed in each of the following cases:
$x=v_{1}$ $\implies v_{1}-v_{3}-v_{8}-v_{9}-v_{1},$
$x=v_{2}$ $\implies v_{2}-v_{3}-v_{8}-v_{9}-v_{2},$
$x=v_{4}$ $\implies v_{4}-v_{7}-v_{8}-v_{9}-v_{4},$
$x=v_{6}$ $\implies v_{6}-v_{7}-v_{8}-v_{9}-v_{6}.$
$\diamond$
Also, since $v_{9}$ is nonadjacent with both $v_{4}$ and $v_{2}$, by $\operatorname{diam}(G)=2$ both $N(v_{9})\cap N(v_{4})$ and $N(v_{9})\cap N(v_{2})$ are non-empty.
Claim 1.6. There is a vertex $v_{10}\in(N(v_{9})\cap N(v_{4}))\cup(N(v_{9})\cap N(v_{2}))$ such that
$v_{10}\notin(\{v_{7},v_{8}\}\cup V(G^{\prime}))$.
Proof. If $v_{8}\in(N(v_{9})\cap N(v_{4}))\cup(N(v_{9})\cap N(v_{2}))$ then $v_{8}$ is adjacent to one of $v_{2}$ or $v_{4}$ and, by Claim 1.4, a 44–cycle is present. Thus, $v_{10}\neq v_{8}$. Now, suppose $v_{10}\in(N(v_{9})\cap N(v_{4}))\cup(N(v_{9})\cap N(v_{2}))$ such that $(N(v_{9})\cap N(v_{4}))\cup(N(v_{9})\cap N(v_{2}))\subseteq(\{v_{7}\}\cup V(G^{\prime}))$. Then, by the diameter 2 condition and Claim 1.5, $v_{9}$ is adjacent to both $v_{3}$ and $v_{7}$ (else $v_{9}$ would be adjacent to another pair of vertices in $\{v_{7}\}\cup V(G^{\prime})$ violating Claim 1.5). However, this forms a 44–cycle by $v_{3}-v_{8}-v_{7}-v_{9}-v_{3}$, providing a contradiction on the claim of no 44–cycles. $\diamond$
Thus, by Claim 1.6, there is a distinct vertex $v_{10}$ that is in at least one of $N(v_{9})\cap N(v_{2})$ or $N(v_{9})\cap N(v_{4})$ such that $v_{10}\notin(\{v_{7},v_{8}\}\cup V(G^{\prime}))$. In either case, an 88–cycle is formed (see Figure 5 and 6) by:
$v_{10}\in N(v_{9})\cap N(v_{2})$ $\implies v_{10}-v_{2}-v_{3}-v_{1}-v_{4}-v_{7}-v_{8}-v_{9}-v_{10},$
$v_{10}\in N(v_{9})\cap N(v_{4})$ $\implies v_{10}-v_{4}-v_{1}-v_{2}-v_{6}-v_{7}-v_{8}-v_{9}-v_{10}.$
$v_{7}$$v_{4}$$v_{1}$$v_{2}$$v_{6}$$v_{3}$$v_{8}$$v_{9}$$v_{10}$
Figure 5: $v_{10}\in N(v_{2})\cap N(v_{9})$ forming an 88-Cycle
$v_{7}$$v_{4}$$v_{1}$$v_{2}$$v_{6}$$v_{3}$$v_{8}$$v_{9}$$v_{10}$
Figure 6: $v_{10}\in N(v_{4})\cap N(v_{9})$ forming an 88-Cycle
Case 2: $v_{3}\neq v_{5}$
Assume $v_{3}\neq v_{5}$.
Claim 2.1. If $xy\in E(G)$ for some $x\in\{v_{3},v_{4}\}$ and $y\in\{v_{5},v_{6}\}$, then $G$ contains a $4$-cycle.
Proof. Since $v_{1}x\in E(G)$ and $v_{2}y\in E(G)$, the cycle $v_{1}-x-y-v_{2}-v_{1}$ has length $4$. $\diamond$
Hence we may assume that no edge joins the sets $\{v_{3},v_{4}\}$ and $\{v_{5},v_{6}\}$.
Claim 2.2. There exist $a\in\{v_{3},v_{4}\}$ and $b\in\{v_{5},v_{6}\}$ such that
$N(a)\cap N(b)\not\subseteq\{v_{1},v_{2}\}.$
Proof. Because $\operatorname{diam}(G)=2$ and $ab\notin E(G)$ (by Claim 2.1), every such pair $(a,b)$ has a common neighbor, so $N(a)\cap N(b)\neq\varnothing$.
Suppose for contradiction that for every $a\in\{v_{3},v_{4}\}$ and $b\in\{v_{5},v_{6}\}$ we have
$N(a)\cap N(b)\subseteq\{v_{1},v_{2}\}.$
Since $v_{1}$ is adjacent to both $v_{3}$ and $v_{4}$, the vertex $v_{1}$ is a common neighbor of $(a,b)$ exactly when $v_{1}b\in E(G)$. Similarly, since $v_{2}$ is adjacent to both $v_{5}$ and $v_{6}$, the vertex $v_{2}$ is a common neighbor of $(a,b)$ exactly when $v_{2}a\in E(G)$. Thus, for each pair $(a,b)$, at least one of the edges $v_{1}b$ or $v_{2}a$ must exist.
If $v_{1}$ is adjacent to both $v_{5}$ and $v_{6}$, then $v_{1}-v_{5}-v_{2}-v_{6}-v_{1}$ is a $4$-cycle. If $v_{2}$ is adjacent to both $v_{3}$ and $v_{4}$, then $v_{1}-v_{3}-v_{2}-v_{4}-v_{1}$ is a $4$-cycle. Hence $v_{1}$ is adjacent to at most one of $\{v_{5},v_{6}\}$ and $v_{2}$ is adjacent to at most one of $\{v_{3},v_{4}\}$. But then there is some pair $(a,b)$ with $v_{1}b\notin E(G)$ and $v_{2}a\notin E(G)$, contradicting the requirement that each pair $(a,b)$ has a common neighbor in $\{v_{1},v_{2}\}$. Therefore the claim holds. $\diamond$
By Claim 2.2, choose $a\in\{v_{3},v_{4}\}$ and $b\in\{v_{5},v_{6}\}$ and a vertex
$v_{7}\in N(a)\cap N(b)\setminus\{v_{1},v_{2}\}.$
If $v_{7}\in\{v_{3},v_{4},v_{5},v_{6}\}$, then a $4$-cycle occurs:
$v_{7}=v_{3}$ $\Rightarrow v_{1}-v_{3}-b-v_{2}-v_{1},$
$v_{7}=v_{4}$ $\Rightarrow v_{1}-v_{4}-b-v_{2}-v_{1},$
$v_{7}=v_{5}$ $\Rightarrow v_{1}-a-v_{5}-v_{2}-v_{1},$
$v_{7}=v_{6}$ $\Rightarrow v_{1}-a-v_{6}-v_{2}-v_{1}.$
Thus,
$v_{7}\notin V(G^{\prime})=\{v_{1},v_{2},v_{3},v_{4},v_{5},v_{6}\}.$
Without loss of generality, let $a=v_{3}$ and $b=v_{5}$; hence
$v_{7}\in N(v_{3})\cap N(v_{5}),\quad v_{7}\notin V(G^{\prime})$
(see Figure 7).
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$
Figure 7: A new common neighbor $v_{7}\in N(v_{3})\cap N(v_{5})$ with $v_{7}\notin V(G^{\prime})$.
Claim 2.3. $v_{7}v_{4}\notin E(G)$.
Proof. If $v_{7}v_{4}\in E(G)$, then $v_{4}-v_{7}-v_{3}-v_{1}-v_{4}$ is a $4$-cycle. $\diamond$
Consider the nonadjacent pair $(v_{4},v_{6})$. If $v_{4}v_{6}\in E(G)$, then Claim 2.1 is violated and there is a $4$-cycle. Hence $v_{4}v_{6}\notin E(G)$ and
$N(v_{4})\cap N(v_{6})\neq\varnothing.$
Choose
$v_{8}\in N(v_{4})\cap N(v_{6}).$
Claim 2.4. $v_{8}\notin\{v_{3},v_{5},v_{7}\}$.
Proof. If $v_{8}=v_{3}$ or $v_{8}=v_{5}$, then Claim 2.1 is violated. If $v_{8}=v_{7}$, then Claim 2.3 is violated. $\diamond$
Subcase 2A: $v_{8}=v_{1}$. Then $v_{1}v_{6}\in E(G)$. Since $\delta(G)\geq 3$, choose
$w\in N(v_{6})\setminus\{v_{1},v_{2}\}.$
By Claim 2.1, $w\notin\{v_{3},v_{4}\}$; moreover $w\neq v_{5}$ or $v_{7}$, else a $4$-cycle occurs. The vertices $v_{4}$ and $w$ are nonadjacent, or else $v_{4}-w-v_{6}-v_{1}-v_{4}$ is a $4$-cycle. Thus $\operatorname{diam}(G)=2$ yields
$z\in N(v_{4})\cap N(w),$
with $z$ distinct from all previously defined vertices. Then
$v_{7}-v_{3}-v_{1}-v_{4}-z-w-v_{6}-v_{2}-v_{5}-v_{7}$
contains a cycle of length $8$.
Subcase 2B: $v_{8}=v_{2}$. This case is handled analogously to $v_{8}=v_{1}$: using $\delta(G)\geq 3$, choose $w\in N(v_{4})\setminus\{v_{1},v_{2}\}$, giving $z\in N(w)\cap N(v_{6})$ via $\operatorname{diam}(G)=2$. Thus, obtaining an $8$-cycle.
Subcase 2C: $v_{8}\notin\{v_{1},v_{2}\}$. Then the edges
$v_{7}v_{3},\ v_{3}v_{1},\ v_{1}v_{4},\ v_{4}v_{8},\ v_{8}v_{6},\ v_{6}v_{2},\ v_{2}v_{5},\ v_{5}v_{7}$
form the simple cycle
$v_{7}-v_{3}-v_{1}-v_{4}-v_{8}-v_{6}-v_{2}-v_{5}-v_{7},$
which has length $8$.
Thus, in all cases, $G$ contains a cycle of length $4$ or $8$ (see Figure 8).
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$
Figure 8: Case 2: if $v_{3}\neq v_{5}$, an 8-cycle appears.
$\square$
Acknowledgments
The author thanks Dr. Michael Albert, Editor-in-Chief of the Australasian Journal of Combinatorics, and one anonymous external expert for their careful reading of an earlier draft, advice, and for comments on the originality and merit of this work. Also, thank you to Dr. Tao Wang of Henan University for reading the first draft and presenting a counterexample to case 1 in personal communication that was accounted for in the author’s original notes but not in the first draft.